# Conditioning LLMs to Generate Code-Switched Text

### v2

@@ Appendix A Prompt-tuning for CS-EN translation
Regarding the few-shot strategies, we find out that giving some examples to the models results in outputs that are more aligned with the expected output, which is logical, since this allows the models to more faithfully replicate the examples provided. The more examples given, the more the model is able to comply to leaving the punctuation marks as they are and not standardizing the spelling, but also it tends to add more colloquial terms and alternate spellings.
@@ Appendix D Fine-tuning / Few-shot prompting
In Table 7, we can see the prompt used for a)fine-tuning of base models; b) fine-tuning of instruction-tuned models; and c) 5-shot prompting. For both fine-tuning settings, at inference time the second part of the prompt that contains the target CS sentence is left blank for the model to complete.
@@ Appendix E Pairwise Annotation Guidelines
In case no criteria is applicable to a pair, we ask the annotators to choose their preferred sentence, using their own judgement o additional criteria they might observe in the specific pair of sentences.
@@ Appendix G Implementation of GPT as judge
For the implementation of GPT as judge, the developer and user prompts in Table 8 have been used to prompt GPT-4o. To calculate the scores, we first check the answers the directly contain the desired format, "A", "B or "T", which are the most common. For the rest of the outputs that did not follow this format, it is possible to extract the labels using simple regular expressions. Then, the scores are calculated just like the human score: every time a model is voted, it gets $1$ point, and the loser gets 00 points; in case of ties, both models get $0.5$ points each.

### v3

@@ Appendix A Prompt-tuning for CS-EN translation
Regarding the few-shot strategies, we find out that giving some examples to the models results in outputs that are more aligned with the expected output, which is logical, since this allows the models to more faithfully replicate the examples provided. The more examples given, the more the model is able to comply to leaving the punctuation marks as they are and not standardizing the spelling, but also it tends to add more colloquial terms and alternate spellings. Here are the final 5 examples that were selected as they cover the majority of the phenomena that models were observed to struggle with:
Input
cuando me gusta algo nunca lo hay mi fucking size o no tengo el dinero.
Output
when I like something there’s never my size or I don’t have the money.
Input
excelente compartir contigo gracias por tu amistad <user> u rock
Output
excellent sharing with you thank you for your friendship <user> u rock
Input
fuhk it tacos de frijol
Output
fuhk it bean tacos
Input
<user> como se llama esa app i wanna play it lmfao
Output
<user> what’s that app called i wanna play it lmfao
Input
i tried putting fake eyelashes on rn lmao me ebarre de glue todo el pinche ojo jajajaja #osoalmil jajaja
Output
i tried putting fake eyelashes on rn lmao i put glue all over my damn eye hahahaha #superclumsy hahaha
@@ Appendix D Fine-tuning / Few-shot prompting
In Table 8, we can see the prompt used for a)fine-tuning of base models; b) fine-tuning of instruction-tuned models; and c) 5-shot prompting. For both fine-tuning settings, at inference time the second part of the prompt that contains the target CS sentence is left blank for the model to complete. For the few-shot prompting approach, the examples are selected randomly among the rest of the instances of the test set.
@@ Appendix E Pairwise Annotation Guidelines
In case no criteria is applicable to a pair, we ask the annotators to choose their preferred sentence, using their own judgment o additional criteria they might observe in the specific pair of sentences.
@@ Appendix G Implementation of GPT as judge
For the implementation of GPT as judge, the developer and user prompts in Table 9 have been used to prompt GPT-4o. To calculate the scores, we first check the answers the directly contain the desired format, "A", "B or "T", which are the most common. For the rest of the outputs that did not follow this format, it is possible to extract the labels using simple regular expressions. Then, the scores are calculated just like the human score: every time a model is voted, it gets $1$ point, and the loser gets 0 points; in case of ties, both models get $0.5$ points each.